OpenStudy (anonymous):

Blake has a total of 64 DVD's and CD's. If the number of CD's is eight more than three times the number of DVD's, how many DVD's does he have?

OpenStudy (anonymous):

Do your own homework!

OpenStudy (farmdawgnation):

Gentlemen, none of these responses are acceptable answers to this question. Knock it off.

OpenStudy (anonymous):

^thanks

OpenStudy (rulnick):

c = 8+3d c+d=64 so c = 50 (CDs) d=14 (DVDs)

OpenStudy (anonymous):

thanks rulnick

OpenStudy (rulnick):

you're welcome

OpenStudy (anonymous):

Anytime babe